OpenStudy (anonymous):

Let f(x) = 2x - 6 Solve f-1(x) when x = 2.

OpenStudy (anonymous):

-2

OpenStudy (austinl):

Do you mean, \(f^{-1}(x)\)

OpenStudy (austinl):

@sabrinasabor111 we don't just give answers here. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

yes @austinL thats what it is

OpenStudy (austinl):

Ok, we have: \(f(x) = 2x - 6\) Do find the inverse, we replace f(x) with y, and then solve for x. Once we have done that, we swap the x and y variables, then change y to f^-1(x). \(f(x) = 2x - 6\) \(y=2x-6\) \(2x=y+6\) \(\displaystyle x=\frac{y+6}{2}\) \(\displaystyle y=\frac{x+6}{2}\) \(\displaystyle f^{-1}(x)=\frac{x+6}{2}\) Then to find the last part, you just input 2 for all instances of x! \(f^{-1}(2)=~?\)

OpenStudy (anonymous):

4

OpenStudy (austinl):

Correctomundo!

OpenStudy (anonymous):

thanks can you help me with another one please?

OpenStudy (austinl):

Sure.

OpenStudy (anonymous):

Let f(x) = 2x + 2. Solve f-1(x) when x = 4.

OpenStudy (austinl):

Ok, how would you find the inverse of that f(x)? Use the exact same steps as above :)

OpenStudy (anonymous):

f(x)= 2x+2 y=2x+2 2x= y+ 2 x= y+2 ---- 2 y=x+2 ----- 2 f^-1(x)=x+2 ----- 2 (4)+2 ----- 2 6 - 2 3 is the final answer

OpenStudy (anonymous):

@austinL did I get it?

OpenStudy (austinl):

Um, back at the top of your work. y=2x+2 What do you do to cancel out an addition?

OpenStudy (anonymous):

-2

OpenStudy (austinl):

Exactly, so we would have, 2x= y - 2

OpenStudy (austinl):

Can you continue from there?

OpenStudy (anonymous):

2x=y-2 x=y-2 ---- 2 y=y-2 --- 2 4-2 --- 2 2 -- 2 so the answer is 1

OpenStudy (austinl):

I'd think so :)

OpenStudy (austinl):

Good job!

OpenStudy (anonymous):

okay I have one last question: Let f(x) = x^2 - 16. Find f-1(x). I don't know how to do it because of the exponent

OpenStudy (austinl):

To cancel out a square, you take the square root :) Make sense?

OpenStudy (anonymous):

a little care to show an example

OpenStudy (anonymous):

would it become y= x + x - 16?

OpenStudy (austinl):

Lets say, \(f(x)=x^2-25\) \(y=x^2-25\) \(y+25=x^2\) \(x=\sqrt{y}+5\) Then swap etc :) Now can you apply it to your problem?

OpenStudy (anonymous):

f(x) = X^2 -16 y = x^2 - 16 y + 16 = x^2 x= \[\sqrt{y+16}\]

OpenStudy (anonymous):

y= \[\sqrt{x+16}\]

OpenStudy (austinl):

\(f^{-1}(x)=\sqrt{x}+4\)

OpenStudy (anonymous):

okay thank you for all your help!

OpenStudy (austinl):

No problem, my pleasure!